The axis of symmetry of a parabola is the line x=4.
The x-intercepts are 10 and -2.
Suppose that the point (12,4) is on the graph;what other point must also be on the graph?
I keep getting the point (4,4) would be on the graph but it is wrong. What am I doing wrong and what point is on the graph? Thanks for any help.

Answer by TimothyLamb(4379)

the axis of symmetry of the parabola is the line x = 4
---
this means that you can fold a graph of the parabola along the above line and the two halves of the graph will overlap perfectly.
---
if the point p1 (12, 4) is on the parabola then another point p2 symmetrical to it must also be on the parabola.
---
reflect p1 (12, 4) about the axis of symmetry x = 4
the other point p2 is at distance dx from the axis of symmetry:
dx = 12 - 4 = 8
but on the opposite side of the axis of symmetry:
---
Answer:
the other point (p2) is:
x = 4 - dx
x = 4 - 8
x = -4
y = 4 (from the y-coordinate of p1)
(-4, 4)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php